# Upper bounds of focusing light through multimode fibers

### v1

@@ Abstract
Wavefront shaping enables precise control of light propagation through multimode fibers, facilitating diffraction-limited focusing for applications such as high-resolution single-fiber imaging and high-power fiber amplifiers. While the theoretical intensity enhancement at the focal point is dictated by the number of input degrees of freedom, practical constraints—such as phase-only modulation and experimental noise—impose significant limitations. Despite its importance, the upper bounds of enhancement under these constraints remain largely unexplored. In this work, we establish a theoretical framework to predict the fundamental limits of intensity enhancement with phase-only modulation in the presence of noise-induced phase errors, and we experimentally demonstrate wavefront shaping that approaches these limits. Our experimental results confirm an enhancement factor of 5,000 in a large-core multimode fiber, approaching the theoretical upper bound, enabled by noise-tolerant wavefront shaping. These findings provide key insights into the limits of phase-only control in multimode fibers, with profound implications for single-fiber imaging, optical communication, high-power broad-area fiber amplification, and beyond.
@@ I Introduction
Multimode fibers (MMFs) are essential for high-resolution and ultra-thin single-fiber endoscopic imaging Čižmár and Dholakia (2012); Choi et al. (2012); Papadopoulos et al. (2012, 2013); Loterie et al. (2015); Amitonova and De Boer (2018); Amitonova and de Boer (2020), optical manipulation Čižmár and Dholakia (2011); Leite et al. (2018), high-bandwidth short-distance optical communication Li et al. (2021a), precise laser-based material processing Kakkava et al. (2019), and power scaling in fiber amplifiers Dawson et al. (2008); Richardson et al. (2010); Zervas and Codemard (2014); Fu et al. (2017). Their ability to support a large number of spatial modes increases the capacity for information transmission and energy delivery but also introduces challenges due to modal dispersion and complex interference effects. Overcoming these challenges requires precise control over light propagation within MMFs. Wavefront shaping provides a powerful approach to manipulate interference at the fiber’s output by tailoring the input field, enabling applications such as high-resolution imaging, targeted light delivery, and nonlinear effect management Di Leonardo and Bianchi (2011); Mosk et al. (2012); Horstmeyer et al. (2015); Vellekoop (2015); Rotter and Gigan (2017); Tzang et al. (2018); Gigan et al. (2022); Cao et al. (2022, 2023); Chen et al. (2023); Wisal et al. (2024a). Precise control of the output field through input wavefront shaping requires accurately measuring the fiber’s transmission matrix, which characterizes the input-output field relationship Popoff et al. (2010); Čižmár and Dholakia (2011); Choi et al. (2012); Kim et al. (2015); Li et al. (2021b); Bender et al. (2023). This measurement is essential for optimizing wavefront shaping techniques, including the fundamental task of focusing light at a desired location Vellekoop and Mosk (2007); Vellekoop et al. (2010); Papadopoulos et al. (2012); Caravaca-Aguirre et al. (2013); Amitonova et al. (2016); Descloux et al. (2016); Lyu et al. (2022); Gomes et al. (2022); Lyu and Amitonova (2024).
@@
A key metric for evaluating wavefront shaping performance for focusing light through complex media is the enhancement factor, which quantifies the intensity at the focal position Vellekoop and Mosk (2007); Amitonova et al. (2016). The enhancement factor is defined as $\eta_{m}=I_{m}^{(\mathrm{foc})}/\langle I_{m}^{(\mathrm{rand})}\rangle$, where $I_{m}^{(\mathrm{foc})}$ is the focus intensity at the target position $m$ after wavefront shaping and $\langle I_{m}^{(\mathrm{rand})}\rangle$ is the averaged intensity at the same position $m$ over multiple random wavefront inputs Vellekoop and Mosk (2007); Vellekoop (2015); Mastiani et al. (2024). To maximize the enhancement factor, an optimization algorithm is typically used to determine the ideal input wavefront Vellekoop and Mosk (2008a); Yılmaz et al. (2013); Mastiani and Vellekoop (2021); Mastiani et al. (2024). Fundamentally, these algorithms are based on some form of transmission matrix measurement. In principle, with complete control over both the amplitude and phase of the input channels, all transmitted power can be efficiently collected at a desired focal position Gomes et al. (2022). However, in practice, there are certain limitations to the focus intensity due to incomplete input channel control Resisi et al. (2020); Mastiani et al. (2024). With phase and amplitude modulation of the input wavefront, the enhancement factor is equal to the number of input degrees of freedom denoted as $\eta=N$. Nevertheless, most spatial light modulators (SLM) function primarily as phase-only modulators, as amplitude modulation is generally avoided in practice Mosk et al. (2012). This is because amplitude modulation reduces the input power, making it less efficient for applications requiring high-intensity light focusing, such as broad-area fiber amplifiers and laser ablation through large-core optical fibers Kakkava et al. (2019); Wisal et al. (2024b); Rothe et al. (2025a, b). Therefore, in practical scenarios, phase-only input modulation is typically preferred; the input wavefronts can be optimally shaped to achieve constructive interference at the desired focus location at the fiber’s output, maintaining a constant input power Gigan et al. (2022); Cao et al. (2023); Vellekoop and Mosk (2008a); Rothe et al. (2025a). This constraint inherently limits the maximum enhancement factor, with its theoretical upper bound, following $\eta=R(N-1)+1$, where $R=\pi/4$ represents the participation ratio—a well-established result for focusing light through disordered scattering media. Vellekoop and Mosk (2007); van Beijnum (2009); Vellekoop and Mosk (2008a); Yılmaz et al. (2013); Vellekoop (2015); Mastiani et al. (2024).
@@
Here, we present a comprehensive work that combines both experimental and theoretical aspects of wavefront shaping. We introduce the theoretical upper bound on the enhancement factor for phase-only modulation; moreover, we experimentally demonstrate it for focusing light through a multimode fiber (MMF). We first describe how the participation ratio $R$ depends on the basis used for phase-only modulation. Our theoretical calculations and experimental observations reveal that when wavefront modulation is performed on the Fourier space of the proximal end of the fiber, the $R_{m}=\pi/4$ at any fiber output position $m$, which closely resembles the well-known result observed in disordered media. In contrast, phase-only modulation in the fiber mode basis reveals a strong radial dependence of $R$ at the fiber’s distal end. To experimentally approach the theoretical upper bound of the enhancement factor for phase-only modulation, we perform noise-tolerant transmission matrix measurements using a Hadamard basis, achieving an enhancement factor of $\eta=5,000$. We also introduce a predictive method that combines our theoretical framework with measured phase errors in a practical wavefront shaping setup. Using this approach, we quantify the phase errors in transmission matrix measurements both in the canonical (SLM pixel) and Hadamard bases. This method provides a clear quantitative understanding of how Hadamard-based measurements minimize phase errors, enabling near-ideal wavefront shaping. By linking theoretical predictions with experimental observations, this work establishes a foundation for accurately predicting the enhancement factor and achieving near-perfect phase-only wavefront shaping in multimode fibers. Our findings directly advance the development of more robust and efficient wavefront shaping techniques, with broad applications in high-resolution imaging, broad-area fiber amplifiers, laser ablation through multimode fibers, and beyond.
@@
II Transmission matrix and focusing measurements
@@
Using the spatial light modulator in our experimental setup, shown in Fig. 1(a), we select an input basis indexed by $n$ (either canonical or Hadamard) for the incident fields $E_{n}$, as defined in Equation 1. The field transmission matrix is then measured using common-path phase-shifting interferometry Popoff et al. (2010); Hsu et al. (2017); Yılmaz et al. (2019, 2019) in a chosen basis.
@@
Figure 2: (a) The mean participation ratio $R$, averaged over the core radius $a$ and the number of degrees of freedom $N$, versus the normalized radial distance $r/a$. The experimental $R$ (solid red line) is consistent with the numerical $R$ (solid blue line) in the Fourier basis, both following a trend close to $\pi/4$ (dashed black line) and showing no dependence on the radial distance. However, we observe a strong dependence of $R$ on the radial distance when the $R$ is computed with phase-only modulation on the MMF fiber mode basis (solid black line). (b) The mean participation ratio $R$, averaged over the core radius $a$ and the radial distance $r$ at the fiber distal end, shown with respect to the normalized number of degrees of freedom $N/N_{\text{mode}}$. The experimental $R$ (red line) agrees with the numerical $R$ (solid blue line) and remains invariant with $N/N_{\text{mode}}$, maintaining a value near $\pi/4$ (dashed black line). (c) The mean enhancement factor $\eta$, averaged over the azimuthal position $\theta$ and radial distances $r$ at the fiber distal end, shown with respect to the number of degrees of freedom $N$ = 172, 484, 952, and 2,032. The experimental enhancement factor $\eta$ in the Hadamard basis (red solid line) closely follows the theoretical prediction (black solid line) and is notably higher than the experimental $\eta$ in the canonical basis (blue solid line).
@@
To measure the transmission matrix of the MMF using common-path interferometry Popoff et al. (2010), the light field on the spatial light modulator is divided into a signal and a reference part. The signal is modulated with four phase steps uniformly spaced between 00 and $2\pi$, and both the signal and reference components are then coupled into the MMF. The MMF is a step-index fiber with a diameter of 200 µm, a length of 6 cm, and an NA = 0.22. The light propagates through the supported modes of the MMF, and the resulting intensity pattern at the distal end is imaged onto a charged-coupled device (CCD) camera sensor. In this work, we use feedback-based wavefront shaping to retrieve the transmission matrix from the SLM (in the Fourier space) to the CCD camera (in the real space) using two different algorithms: the stepwise sequential algorithm (SSA) on the canonical basis Vellekoop and Mosk (2007) and the dual reference algorithm on the Hadamard basis Mastiani and Vellekoop (2021). For both methods, we divide the SLM surface into $N$ number of input degrees of freedom (segments) where $N$ is chosen to be 172, 484, 952, 2,032, 3,300, 3,940, 5,388, 6,180, 7,080, and 8,000 in all the experiments described here. We choose $N$ in such a way that we display a circular phase pattern on the SLM. In our experimental setup, depicted in Fig. 1(a), we modulate a single linear horizontal polarization of light at the input and detect the same linear horizontal polarization at the output (the details are described in the supplementary material, section I - A).
@@
Maximizing the overlap between the input field patterns and the fiber’s core at its proximal end, both in real and Fourier space, is essential. In real space, the input field patterns must be centered at the fiber core, and their size should be equal to or smaller than the fiber core diameter. In Fourier space, the diameter of the displayed SLM pattern should correspond to a numerical aperture equal to or smaller than the fiber’s numerical aperture, NA = 0.22. We adjusted our setup such that displaying SLM patterns with $6\times 6$ pixel-size segments ensures a near-perfect size match between any arbitrary SLM pattern and the fiber core (additional details can be found in the supplementary material, section I - A). Furthermore, the diameter of the SLM pattern consistently corresponds to an NA smaller than that of the fiber, with NA $<$ 0.22.
@@
The phases of each transmission matrix column $m$ are computed using four recorded CCD camera images, each corresponding to a different relative phase between the signal and the reference, obtained through common-path four-phase shifting interferometry with the SLM Popoff et al. (2010); Hsu et al. (2017); Yılmaz et al. (2019, 2019). The transmission matrix elements map the input field onto the output field, where every element on the output field is a summation of all the input elements multiplied by an analogous transmission coefficient. Thus, a random summation of $N$ field components contributes to each element in the output field, which results in the speckle pattern as is seen in Fig. 1(b). To form a focus, i.e., to increase light intensity for a specific output element, the $N$ field components must constructively interfere with each other. This is accomplished by displaying the conjugate of the measured transmission matrix phase row with index $m$, which corresponds to the desired position of the output field, on the SLM.
@@
In the dual reference algorithm Mastiani and Vellekoop (2021), we use the Hadamard basis, where the size of the basis must be $N_{1}=2^{p}$, with $p$ being an integer (e.g., $N_{1}$ = 128, 256, 512, 1,024, 2,048, 4,096). In this approach, we divide the SLM segments into two equal-sized groups, each containing a small number of overlapping segments, denoted as $O$. In the first step of the algorithm, a Hadamard pattern is displayed on the segments of group 1 (segments from 1 to $N_{1}$), while the remaining segments (from $N_{1}+1$ to $N$) contribute as the reference field $E_{\rm ref}^{(1)}$. Four-phase shift interferometry is applied to both groups to vary their relative phase from 0 to $2\pi$ Popoff et al. (2010). Subsequently, we perform a Hadamard transform on the full-field patterns obtained from the four-phase-shifting interferometry to extract the transmission matrix phase elements for group 1 relative to the $E_{\rm ref}^{(1)}$ phase, which results in the transmission matrix 1 $t^{(1)}$.
@@
Note that, in all transmission matrix experiments, we recorded fiber output intensity images $I_{m}^{(\mathrm{rand})}$ for 1,000 random wavefront inputs, using the same number of segments $N$ as in each experiment. Fig. 1(b) shows an example random speckle pattern observed on the CCD camera for a random input wavefront. These speckles result from modal dispersion, where each fiber mode propagates with a distinct propagation constant and phase delay, leading to complex interference at the output.
Immediately after each transmission matrix measurement, we perform focusing experiments. To focus at a specific position $m$ within the core at the fiber’s distal end, we display the conjugated phase from the measured transmission matrix row corresponding to position $m$ and record the focused intensity pattern on the CCD camera. To avoid CCD camera saturation, calibrated neutral density (ND) filters were placed in front of the camera during the measurements. These experiments are performed to focus on various radial $(r)$ and azimuthal $(\theta)$ positions within the core of the fiber’s distal end using phase-only input modulation, obtained from the corresponding transmission matrix data. The resulting enhancement factors are calculated using the expression $\eta_{m}=I_{m}^{(\mathrm{foc})}/\langle I_{m}^{(\mathrm{rand})}\rangle$, where $I_{m}^{(\mathrm{foc})}$ represents the focused intensity at the target position $m$ after wavefront shaping, and $\langle I_{m}^{(\mathrm{rand})}\rangle$ represents the averaged intensity at the same position $m$ over 1,000 random wavefront inputs. Two example intensity patterns are shown in Fig. 1(c) and Fig. 1(d) for wavefront shaping on the canonical and Hadamard basis, respectively. As evident from the captured images, the optimized focus achieved using the Hadamard basis exhibits a higher peak intensity compared to the focus obtained on the canonical basis.
@@
Figure 3: Phase error measurements in wavefront shaping experiments. (a) A conceptual sketch of two independently measured phase maps, $\phi_{mn}^{(1)}$ and $\phi_{mn}^{(2)}$ for the same input $n$. The phase difference, $\delta\phi_{mn}^{\prime}=\arg\left(e^{i(\phi_{mn}^{(2)}-\phi_{mn}^{(1)})}\right)$, represents the measured phase error. (b) Standard deviation $\sigma\mathrm{{}_{\delta\phi}}$ of the Gaussian-fitted phase error distributions as a function of the number of degrees of freedom $N$. In the canonical basis, $\sigma\mathrm{{}_{\delta\phi}}$ increases with $N$, indicating higher phase errors. In contrast, the Hadamard basis maintains a consistently low $\sigma\mathrm{{}_{\delta\phi}}$, suggesting greater robustness to phase errors. (c, d) Probability density functions of the phase errors $P(\delta\phi_{mn}^{\prime})$ for the canonical and Hadamard bases, respectively, as a function of $N$. The Gaussian-fitted curves (red) show that in the canonical basis, the phase error distribution broadens significantly with increasing $N$. In contrast, (d) shows that the Hadamard basis maintains a sharply peaked distribution around zero, indicating minimal phase errors.
Figure 4: The impact of phase errors on the enhancement factor. (a) Schematic representation of the phase error coefficient $\Phi$ for the Hadamard (red) and canonical (blue) bases, where $\delta\phi$ represents phase errors. The dashed lines indicate how phase errors accumulate differently in the two bases, with the Hadamard basis maintaining $\Phi\approx 1$ due to a reduced cumulative effect of phase errors. (b) The phase error coefficient $\Phi$ as a function of the number of degrees of freedom $N$. On the canonical basis, $\Phi$ decreases with $N$ due to a decreasing signal-to-noise ratio (SNR), which reduces interferometric visibility. In contrast, the Hadamard basis maintains $\Phi\approx 1$ due to a balanced signal-to-background ratio. (c) The phase error coefficient $\Phi$ for $N=8000$, plotted as a function of the normalized photon budget. The Hadamard basis maintains a consistently high $\Phi$ compared to the canonical basis, reflecting its improved SNR under the same noise conditions. The normalized photon budget (ranging from 0 to 1) is defined so that a value of 1 corresponds to a maximum photon budget of approximately 214 mean counts on the CCD.

### v3

# Upper bounds on focusing light through multimode fibers

@@ Abstract
Wavefront shaping enables precise control of light propagation through multimode fibers (MMFs), facilitating diffraction-limited focusing for applications such as high-resolution single-fiber imaging and high-power fiber amplifiers. While the theoretical intensity enhancement at the focal point is dictated by the number of input degrees of freedom, practical constraints—such as phase-only modulation and experimental noise—impose significant limitations. Despite its importance, the upper bounds of enhancement under these constraints remain largely unexplored. In this work, we establish a theoretical framework to predict the fundamental limits of intensity enhancement with phase-only modulation in the presence of noise-induced phase errors, and we experimentally demonstrate wavefront shaping that approaches these limits. Our experimental results confirm an enhancement factor of 5000 in a large-core MMF, approaching the theoretical upper bound, enabled by noise-tolerant wavefront shaping. These findings provide key insights into the limits of phase-only control in MMFs, with profound implications for single-fiber imaging, optical communication, high-power broad-area fiber amplification, and beyond.
@@ I Introduction
Multimode fibers (MMFs) are essential for high-resolution and ultra-thin single-fiber endoscopic imaging [1, 2, 3, 4, 5, 6, 7], optical manipulation [8, 9], high-bandwidth short-distance optical communication [10], precise laser-based material processing [11], and power scaling in fiber amplifiers [12, 13, 14, 15]. Their ability to support a large number of spatial modes increases the capacity for information transmission and energy delivery, but also introduces challenges due to modal dispersion and complex interference effects. Overcoming these challenges requires precise control over light propagation within MMFs. Wavefront shaping provides a powerful approach to manipulate interference at the fiber’s output by tailoring the input field using a spatial light modulator (SLM), enabling applications such as high-resolution imaging, targeted light delivery, and nonlinear effect management [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Precise control of the output field through input wavefront shaping requires accurately measuring the fiber’s transmission matrix, which characterizes the input-output field relationship [27, 8, 2, 28, 29, 30]. This measurement is essential for optimizing wavefront shaping techniques, including the fundamental task of focusing light at a desired location [31, 32, 3, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42].
@@
A key metric for evaluating wavefront shaping performance for focusing light through complex media is the enhancement factor, which quantifies the intensity at the focal position [31, 35]. The enhancement factor is defined as $\eta_{m}=I_{m}^{(\mathrm{foc})}/\langle I_{m}^{(\mathrm{rand})}\rangle$, where $I_{m}^{(\mathrm{foc})}$ is the focus intensity at the target position $m$ after wavefront shaping and $\langle I_{m}^{(\mathrm{rand})}\rangle$ is the averaged intensity at the same position $m$ over multiple independent random wavefront inputs [31, 19, 43]. To maximize the enhancement factor, an optimization algorithm is typically used to determine the ideal input wavefront [44, 45, 34, 37, 38, 46, 43]. Fundamentally, these algorithms are based on some form of transmission-matrix measurement. In principle, with complete control over both the amplitude and phase of the input channels, all transmitted power can be efficiently collected at a desired focal position [40]. However, in practice, there are certain limitations to the focus intensity due to incomplete input channel control [47, 43]. With full-field modulation of the input wavefront, the upper bound of the enhancement factor equals the number of controlled channels, $\eta=N$. This result relies on ideal amplitude and phase control, as well as ideal mode coupling—conditions typically satisfied in short, low-loss multimode fibers where mode-dependent loss is negligible (see Appendix B). Nevertheless, most SLMs function primarily as phase-only modulators, as amplitude modulation is generally avoided in practice [17]. This is because amplitude modulation reduces the input power, making it less efficient for applications requiring high-intensity light focusing, such as broad-area fiber amplifiers and laser ablation through large-core optical fibers [11, 48, 49, 50]. Therefore, in practical scenarios, phase-only input modulation is typically preferred; the input wavefronts can be optimally shaped to achieve constructive interference at the desired focus location at the fiber’s output, maintaining a constant input power [22, 24, 44, 49]. This constraint inherently limits the maximum enhancement factor, with its theoretical upper bound, following $\eta=R(N-1)+1$, where $R=\pi/4$ represents the participation ratio—a well-established result for focusing light through disordered scattering media. [31, 51, 44, 45, 19, 43].
@@
Here, we present a comprehensive work that combines both experimental and theoretical aspects of wavefront shaping. We introduce the theoretical upper bound on the enhancement factor for phase-only modulation; moreover, we experimentally demonstrate it for focusing light through a multimode fiber (MMF). We first describe how the participation ratio $R$ depends on the basis used for phase-only modulation. Our theoretical calculations and experimental observations reveal that when wavefront modulation is performed on the Fourier space of the proximal end of the fiber, the $R_{m}=\pi/4$ at any fiber output position $m$, which closely resembles the well-known result observed in disordered media. In contrast, phase-only modulation in the fiber mode basis reveals a strong radial dependence of $R$ at the fiber’s distal end. To experimentally approach the theoretical upper bound of the enhancement factor for phase-only modulation, we perform noise-tolerant transmission-matrix measurements using a Hadamard basis, achieving an enhancement factor of $\eta=5,000$. We also introduce a predictive method that combines our theoretical framework with measured phase errors in a practical wavefront shaping setup. Using this approach, we quantify the phase errors in transmission-matrix measurements both in the canonical (SLM pixel) and Hadamard bases. This method provides a clear quantitative understanding of how Hadamard-based measurements minimize phase errors, enabling near-ideal wavefront shaping. By linking theoretical predictions with experimental observations, this work establishes a foundation for accurately predicting the enhancement factor and achieving near-perfect phase-only wavefront shaping in multimode fibers. Our findings directly advance the development of more robust and efficient wavefront shaping techniques, with broad applications in high-resolution imaging, broad-area fiber amplifiers, laser ablation through multimode fibers, and beyond.
@@
II Transmission-matrix and focusing measurements
@@
Using the spatial light modulator in our experimental setup, shown in Fig. 1(a), we select an input basis indexed by $n$ (either canonical or Hadamard) for the incident fields $E_{n}$, as defined in Equation 1. The field transmission matrix is then measured using common-path phase-shifting interferometry [27, 52, 53, 54] in a chosen basis.
@@
Figure 2: (a) The mean participation ratio $R$, averaged over the core radius $a$ and the number of degrees of freedom $N$, versus the normalized radial distance $r/a$ is illustrated. The experimental $R$ (solid red line) is consistent with the numerical $R$ (solid blue line) in the Fourier basis, both following a trend close to $\pi/4$ (dashed black line) and showing no dependence on the radial distance. However, we observe a strong dependence of $R$ on the radial distance when $R$ is computed with phase-only modulation on the MMF fiber mode basis (solid black line). (b) The mean participation ratio $R$, averaged over the core radius $a$ and the radial distance $r$ at the fiber distal end, is shown with respect to the normalized number of degrees of freedom $N/M$. The experimental $R$ (red line) agrees with the numerical $R$ (solid blue line) and remains invariant with $N/M$, maintaining a value near $\pi/4$ (dashed black line). (c) The mean enhancement factor $\eta$, averaged over the azimuthal position $\theta$ and radial distances $r$ at the fiber distal end, is shown with respect to the number of degrees of freedom NN = 172, 484, 952, and 2,032. The experimental enhancement factor $\eta$ in the Hadamard basis (red solid line) closely follows the theoretical prediction (black solid line) and is notably higher than the experimental $\eta$ in the canonical basis (blue solid line).
@@
To measure the transmission matrix of the MMF using common-path interferometry [27], the light field on the spatial light modulator is divided into a signal and a reference part. The signal is modulated with four phase steps uniformly spaced between 0 and $2\pi$, and both the signal and reference components are then coupled into the MMF. The MMF is a step-index (SI) fiber with a diameter of 200 µm, a length of 6 cm, and an NA = 0.22. The light propagates through the supported modes of the MMF, and the resulting intensity pattern at the distal end is imaged onto a charged-coupled device (CCD) camera sensor. In this work, we use feedback-based wavefront shaping to retrieve the transmission matrix from the SLM (in the Fourier space) to the CCD camera (in the real space) using two different algorithms: the stepwise sequential algorithm (SSA) on the canonical basis [31] and the dual reference algorithm on the Hadamard basis [46]. For both methods, we divide the SLM surface into $N$ number of input degrees of freedom (segments) where $N$ is chosen to be 172, 484, 952, 2,032, 3,300, 3,940, 5,388, 6,180, 7,080, and 8,000 in all the experiments described here. We choose $N$ in such a way that we display a circular phase pattern on the SLM. In our experimental setup, depicted in Fig. 1(a), we modulate a single linear horizontal polarization of light at the input and detect the same linear horizontal polarization at the output (the details are described in Appendix A).
@@
Maximizing the overlap between the input field patterns and the fiber’s core at its proximal end, both in real and Fourier space, is essential. In real space, the input field patterns must be centered at the fiber core, and their size should be equal to or smaller than the fiber core diameter. In Fourier space, the diameter of the displayed SLM pattern should correspond to a numerical aperture equal to or smaller than the fiber’s numerical aperture, NA = 0.22. We adjusted our setup such that displaying SLM patterns with $6\times 6$ pixel-size segments ensures a near-perfect size match between any arbitrary SLM pattern and the fiber core (additional details can be found in Appendix A). Furthermore, the diameter of the SLM pattern consistently corresponds to an NA smaller than that of the fiber, with NA << 0.22.
@@
The phases of each transmission-matrix column is computed using four recorded CCD camera images, each corresponding to a different relative phase between the signal and the reference, obtained through common-path four-phase shifting interferometry with the SLM [27, 52, 53, 54]. The transmission-matrix elements map the input field onto the output field, where every element on the output field is a summation of all the input elements multiplied by an analogous transmission coefficient. Thus, a random summation of $N$ field components contributes to each element in the output field, which results in the speckle pattern as is seen in Fig. 1(b). To form a focus, i.e., to increase light intensity for a specific output element, the $N$ field components must constructively interfere with each other. This is accomplished by displaying the conjugate of the measured transmission-matrix phase row with index $m$, which corresponds to the desired position of the output field, on the SLM.
@@
In the dual reference algorithm [46], we use the Hadamard basis, where the size of the basis must be $N_{1}=2^{p}$, with $p$ being an integer (e.g., $N_{1}$ = 128, 256, 512, 1,024, 2,048, 4,096). In this approach, we divide the SLM segments into two equal-sized groups, each containing a small number of overlapping segments, denoted as $O$. In the first step of the algorithm, a Hadamard pattern is displayed on the segments of group 1 (segments from 1 to $N_{1}$), while the remaining segments (from $N_{1}+1$ to $N$) contribute as the reference field $E_{\rm ref}^{(1)}$. Four-phase shift interferometry is applied to both groups to vary their relative phase from 0 to $2\pi$ [27]. Subsequently, we perform a Hadamard transform on the full-field patterns obtained from the four-phase-shifting interferometry to extract the transmission-matrix phase elements for group 1 relative to the $E_{\rm ref}^{(1)}$ phase, which results in the transmission matrix 1 $t^{(1)}$.
@@
Note that, in all transmission-matrix experiments, we recorded fiber output intensity images $I_{m}^{(\mathrm{rand})}$ for 1,000 independent random wavefront inputs, using the same number of segments $N$ as in each experiment. Fig. 1(b) shows an example random speckle pattern observed on the CCD camera for a random input wavefront. These speckles result from modal dispersion, where each fiber mode propagates with a distinct propagation constant and phase delay, leading to complex interference at the output.
Immediately after each transmission-matrix measurement, we perform focusing experiments. To focus on a specific position $m$ within the core at the fiber’s distal end, we display the conjugated phase from the measured transmission-matrix row corresponding to position $m$ and record the focused intensity pattern on the CCD camera. To avoid CCD camera saturation, calibrated neutral density (ND) filters were placed in front of the camera during the measurements. These experiments are performed to focus on various radial $(r)$ and azimuthal $(\theta)$ positions within the core of the fiber’s distal end using phase-only input modulation, obtained from the corresponding transmission-matrix data. The resulting enhancement factors are calculated using the expression $\eta_{m}=I_{m}^{(\mathrm{foc})}/\langle I_{m}^{(\mathrm{rand})}\rangle$, where $I_{m}^{(\mathrm{foc})}$ represents the focused intensity at the target position $m$ after wavefront shaping, and $\langle I_{m}^{(\mathrm{rand})}\rangle$ represents the averaged intensity at the same position $m$ over 1,000 independent random wavefront inputs. Two example intensity patterns are shown in Fig. 1(c) and Fig. 1(d) for wavefront shaping on the canonical and Hadamard basis, respectively. As evident from the captured images, the optimized focus achieved using the Hadamard basis exhibits a higher peak intensity compared to the focus obtained on the canonical basis.
@@
For imaging applications, not only the enhancement factor but also the spatial shape of the focus, i.e., the point spread function (PSF), is important. While phase-only wavefront shaping in the Fourier plane at the proximal end of the fiber can yield focal spots with a uniform enhancement factor $\eta$ across the core at the distal end, as shown in Fig. 1(e), this does not imply a position-invariant PSF within the core [36].
Figure 3: Phase-error measurements in wavefront shaping experiments are shown. (a) A conceptual sketch of two independently measured phase maps, $\phi_{mn}^{(1)}$ and $\phi_{mn}^{(2)}$ for the same input $n$. The phase difference, $\delta\phi_{mn}^{\prime}=\arg\left(e^{i(\phi_{mn}^{(2)}-\phi_{mn}^{(1)})}\right)$, represents the measured phase error. (b) Standard deviation $\sigma\mathrm{{}_{\delta\phi^{\prime}}}$ of the Gaussian-fitted phase-error distributions as a function of the number of degrees of freedom $N$. In the canonical basis, $\sigma\mathrm{{}_{\delta\phi^{\prime}}}$ increases with $N$, indicating higher phase errors. In contrast, the Hadamard basis maintains a consistently low $\sigma\mathrm{{}_{\delta\phi^{\prime}}}$, suggesting greater robustness to phase errors. (c, d) Probability density functions of the phase errors $P(\delta\phi_{mn}^{\prime})$ for the canonical and Hadamard bases, respectively, as a function of $N$. The Gaussian-fitted curves (red) show that in the canonical basis, the phase-error distribution broadens significantly with increasing NN. In contrast, (d) shows that the Hadamard basis maintains a sharply peaked distribution around zero, indicating minimal phase errors.
Figure 4: The impact of phase errors on the enhancement factor is illustrated. (a) Schematic representation of the phase-error coefficient $\Phi$ for the Hadamard (red) and canonical (blue) bases, where $\delta\phi$ represents phase errors. The dashed lines indicate how phase errors accumulate differently in the two bases, with the Hadamard basis maintaining $\Phi\approx 1$ due to a reduced cumulative effect of phase errors. (b) The phase-error coefficient $\Phi$ as a function of the number of degrees of freedom $N$ is shown. On the canonical basis, $\Phi$ decreases with NN due to a decreasing signal-to-noise ratio (SNR), which reduces interferometric visibility. In contrast, the Hadamard basis maintains $\Phi\approx 1$ due to a balanced signal-to-reference ratio. (c) The phase-error coefficient $\Phi$ for $N=8000$, plotted as a function of the normalized photon budget. The Hadamard basis maintains a consistently high $\Phi$ compared to the canonical basis, reflecting its improved SNR under the same noise conditions. The normalized photon budget (ranging from 0 to 1) is defined so that a value of 1 corresponds to a maximum photon budget of approximately 214 mean counts on the CCD.